# Learning to Partially Defer for Sequences

### v1

@@ Appendix B Proofs for Section 4
${\cal R}_{\text{multi }\psi}(r)=\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)$ ${\cal R}_{\text{multi }\psi}^{*}=\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)$
@@
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)\right)$
@@
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)\right)$
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{\sum_{j\in{\cal J}}A_{j}\psi\left({\boldsymbol{g}}(x),j\right)-\inf_{\tilde{{\boldsymbol{g}}}}\sum_{j\in{\cal J}}A_{j}\psi\left(\tilde{{\boldsymbol{g}}}(x),j\right)}{\Delta}\right)$
@@
=$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\right)$
@@
We know that ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of [55] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
@@
=$\overset{(a)}{=}(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n}\right)-{\cal R}^{*}_{\phi}}{\Delta}\right)$
@@
=$\overset{(b)}{=}0$
(a) by continuity of $\Gamma$ at 00 and (b) since $\Gamma_{\text{ce}}(0)=0$ and $\Gamma_{\text{mae}}(0)=0$
B.5 Proof of Theorem 4.2
@@
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\boldsymbol{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\boldsymbol{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j\in{\cal J}$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$:
@@
=${\cal B}=\{(x,y)\mapsto{\cal L}^{\psi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
@@
$\leq\frac{1}{n}\sup_{r\in{\cal F}}\sum_{j\in{\cal J}}\left|\left(c_{\text{max}}^{i}-l(y^{i},\widehat{y}_{\leq j}^{i})-\tilde{c}_{j}(x^{i},\widehat{y}_{<j}^{i},y^{i})\right)\psi({\boldsymbol{g}}(x^{i}),y^{i})\right|$
$\quad+\left|\left(c_{\text{max}}^{i^{\prime}}-l(y^{i^{\prime}},\widehat{y}_{\leq j}^{i^{\prime}})-\tilde{c}_{j}(x^{i^{\prime}},\widehat{y}_{<j}^{i^{\prime}},y^{i^{\prime}})\right)\psi({\boldsymbol{g}}(x^{i^{\prime}}),y^{i^{\prime}})\right|$
@@
We can further bound $\mathfrak{R}_{n}\left({\cal B}\right)$ by $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ if we show that ${\cal L}^{\psi}$ is Lipschitz continuous with respect to ${\boldsymbol{g}}$
@@
$=\left\|\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}^{1}(x),j)-\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\boldsymbol{g}}^{1}(x),j)-\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
=$\overset{(a)}{=}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\boldsymbol{g}}^{1}(x),j)-\psi({\boldsymbol{g}}^{2}(x),j)\right\|_{2}$
$\overset{(b)}{\leq}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(\bar{C}+\bar{l}\right)\rho\left\|{\boldsymbol{g}}^{1}(x)-{\boldsymbol{g}}^{2}(x)\right\|_{2}$
$=({\cal J}-1)\left(\bar{C}+\bar{l}\right)\rho\left\|{\boldsymbol{g}}^{1}(x)-{\boldsymbol{g}}^{2}(x)\right\|_{2}$
(a) where $j_{\text{max}}=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$, (b) by lipschitz continuity of $\psi$.
By the extension of the Talagrand’s Contraction Lemma [54],
@@
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\rho\sqrt{2}\mathop{\mathbb{E}}_{\sigma,\{z^{i}\}_{i=1}^{n}\sim{\cal P}_{{\cal X}\times{\cal Y}}}\left[\sup_{{\boldsymbol{g}}\in{\cal G}_{1}\times...\times{\cal G}_{L+1}}\frac{1}{n}\sum_{i=1}^{n}\sum_{j\in{\cal J}}\sigma_{i,j}g_{j}(z^{i})\right]$
@@
Suppose $r_{n}\in{\cal F}=\{x\to\mathop{\mathrm{argmin}}\limits_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of $(-M,M)$ with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\psi_{\text{ce}}$ or $\psi_{\text{mae}}$, then with probability $1-\delta$, for a fixed $h$, the excess risk of with respect to ${\cal L}^{\text{OneTime}}$ holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$
@@
where $\tilde{\Gamma}:\mathbb{R}^{+}\to\mathbb{R}^{+}$, $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$, $\rho$ is the lipschitz norm of $\psi$, and ${\cal A}_{\psi}\left({\cal F},h\right)={\cal R}^{*}_{\psi}\left({\cal F},h\right)-{\cal R}^{*}_{\psi}(h)$ is the approximation error.
@@
We know that $\psi_{\text{ce}},\psi_{\text{mae}}$ are upper bounded by $u_{\psi_{\text{ce}}}(M),u_{\psi_{\text{mae}}}(M)$ repsectively. We also know that the lipschitz norm for $\psi_{\text{ce}}$ $\rho_{\text{ce}}$ is $\sqrt{2}$ and the lipschitz norm for $\psi_{\text{mae}}$ $\rho_{\text{mae}}$ is $2$
@@
${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of [55] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
@@ Appendix C Baseline Details
Let $p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ be the predictive distribution of the $j^{\text{th}}$ token conditioned on the leftward context $\widehat{y}_{<j}$ and input $x$. The predictive entropy would then be $H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)=-\sum_{v\in{\cal V}}p_{h}(v\mid\widehat{y}_{<j},x)\log p_{h}(v\mid\widehat{y}_{<j},x)$. For a given threshold $\tau$, the whole deferral decision rule is determined by evaluating $r(x)>\tau$. The baselines offer various ways of modeling $r$.
  • ChowSum - $r_{\text{ChowSum}}(x)=-\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • ChowMean - $r_{\text{ChowMean}}(x)=-\frac{1}{L}\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • ChowQuantile at the level $\alpha$ $r_{\text{ChowQuantile}}^{\alpha}(x)=\text{Quantile}_{\alpha}\left(\left\{-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)\right\}_{j=1}^{L}\right)$
  • WholeModelEmbed - $r(x)$ is trained using the surrogate of the modified One-time deferral loss or ${\cal L}^{\psi}_{\text{mod}}$ where ${\cal J}=\{1,L+1\}$. This is quite similar to the approach proposed by Gupta et al. [40] except we train the classifier with a cost-sensitive loss
Similarly for token-wise deferral, if $r_{j}(x,\widehat{y}_{<j})$ exceeds threshold $\tau$, the expert is called to predict the next token. $r_{j}$ can be modeling in the following ways:
  • TokenwiseSoftmax: $r_{j}^{\text{softmax}}(x,\widehat{y}_{<j})=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • TokenwiseEntropy: $r_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$
Since $r(x)=\mathop{\mathrm{argmax}}\limits_{j\in{\cal J}}g_{j}(x)$ for one-time deferral, we consider the following the scoring functions $g_{j}$:
  • OneTimeSoftmax: $g_{j}^{\text{softmax}}(x)=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$
  • OneTimeEntropy: $g_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$

### v2

@@ Appendix B Proofs for Section 4
${\cal R}_{\text{multi }\psi}(r)=\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\bm{g}}(x),j\right)$ ${\cal R}_{\text{multi }\psi}^{*}=\inf_{\tilde{{\bm{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\bm{g}}(x),j\right)$
@@
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\bm{g}}(x),j\right)-\inf_{\tilde{{\bm{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\bm{g}}}(x),j\right)\right)$
@@
$\leq\sum_{k\in{\cal J}}A_{k}\Gamma\left(\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left({\bm{g}}(x),j\right)-\inf_{\tilde{{\bm{g}}}}\sum_{j\in{\cal J}}\frac{A_{j}}{\sum_{k\in{\cal J}}A_{k}}\psi\left(\tilde{{\bm{g}}}(x),j\right)\right)$
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\frac{\sum_{j\in{\cal J}}A_{j}\psi\left({\bm{g}}(x),j\right)-\inf_{\tilde{{\bm{g}}}}\sum_{j\in{\cal J}}A_{j}\psi\left(\tilde{{\bm{g}}}(x),j\right)}{\Delta}\right)$
@@
$\overset{(b)}{=}\tilde{\Gamma}\left({\cal R}_{\psi}\left(r,h\right)-{\cal R}^{*}_{\psi}(h)\right)$
@@
We know that ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of Mao et al. [2023] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
@@
$\overset{(a)}{=}(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\Gamma\left(\lim_{n\to\infty}\frac{{\cal R}_{\phi}\left(r_{n}\right)-{\cal R}^{*}_{\phi}}{\Delta}\right)$
@@
$\overset{(b)}{=}0$
(a) by continuity of $\Gamma$ at 0 and (b) since $\Gamma_{\text{ce}}(0)=0$ and $\Gamma_{\text{mae}}(0)=0$
B.4 Proof of Theorem 4.2
@@
Suppose $r_{n}\in{\cal F}=\{x\to\operatorname*{arg\,min}_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of (M,M)(-M,M) with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\rho$-lipschitz continuous with respect to ${\bm{g}}(x)$ and there exists $u_{\psi}(M)$ such that $\psi\left({\bm{g}}(x),j\right)\leq u_{\psi}(M)<\infty$ for all $j\in{\cal J}$ and $r\in{\cal F}$, then with probability $1-\delta$ for a fixed $h$, the following upper bound holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$:
@@
${\cal B}=\{(x,y)\mapsto{\cal L}^{\psi}\left(h,r,x,y\right)\ :r\in{\cal F}\}$.
@@
$\leq\frac{1}{n}\sup_{r\in{\cal F}}\sum_{j\in{\cal J}}\left|\left(c_{\text{max}}^{i}-l(y^{i},\widehat{y}_{\leq j}^{i})-\tilde{c}_{j}(x^{i},\widehat{y}_{<j}^{i},y^{i})\right)\psi({\bm{g}}(x^{i}),y^{i})\right|$
$\quad+\left|\left(c_{\text{max}}^{i^{\prime}}-l(y^{i^{\prime}},\widehat{y}_{\leq j}^{i^{\prime}})-\tilde{c}_{j}(x^{i^{\prime}},\widehat{y}_{<j}^{i^{\prime}},y^{i^{\prime}})\right)\psi({\bm{g}}(x^{i^{\prime}}),y^{i^{\prime}})\right|$
@@
We can further bound $\mathfrak{R}_{n}\left({\cal B}\right)$ by $\mathfrak{R}_{n}\left({\cal F}_{j}\right)$ if we show that ${\cal L}^{\psi}$ is Lipschitz continuous with respect to $\bm{g}$}
@@
$=\left\|\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\bm{g}}^{1}(x),j)-\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\psi({\bm{g}}^{2}(x),j)\right\|_{2}$
$\leq\sum_{j\in{\cal J}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\bm{g}}^{1}(x),j)-\psi({\bm{g}}^{2}(x),j)\right\|_{2}$
$\overset{(a)}{=}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(c_{\text{max}}-l(y,\widehat{y}_{\leq j})-\tilde{c}_{j}(x,\widehat{y}_{<j},y)\right)\left\|\psi({\bm{g}}^{1}(x),j)-\psi({\bm{g}}^{2}(x),j)\right\|_{2}$
$\overset{(b)}{\leq}\sum_{j\in{\cal J},j\neq j_{\text{max}}}\left(\bar{C}+\bar{l}\right)\rho\left\|{\bm{g}}^{1}(x)-{\bm{g}}^{2}(x)\right\|_{2}$
$=({\cal J}-1)\left(\bar{C}+\bar{l}\right)\rho\left\|{\bm{g}}^{1}(x)-{\bm{g}}^{2}(x)\right\|_{2}$
(a) where $j_{\text{max}}=\operatorname*{arg\,max}_{j\in{\cal J}}l(y,\widehat{y}_{\leq j})+\tilde{c}_{j}(x,\widehat{y}_{<j},y)$, (b) by lipschitz continuity of $\psi$.
By the extension of the Talagrand’s Contraction Lemma Maurer [2016],
@@
$\leq(|{\cal J}|-1)\left(\bar{C}+\bar{l}\right)\rho\sqrt{2}\mathop{\mathbb{E}}_{\sigma,\{z^{i}\}_{i=1}^{n}\sim{\cal P}_{{\cal X}\times{\cal Y}}}\left[\sup_{{\bm{g}}\in{\cal G}_{1}\times...\times{\cal G}_{L+1}}\frac{1}{n}\sum_{i=1}^{n}\sum_{j\in{\cal J}}\sigma_{i,j}g_{j}(z^{i})\right]$
@@
Suppose $r_{n}\in{\cal F}=\{x\to\operatorname*{arg\,min}_{j\in{\cal J}}g_{j}(x):g_{j}\in{\cal G}_{j}\}$ where ${\cal G}_{j}$ consists of functions having a range of (M,M)(-M,M) with $M\geq 0$, $\tilde{c}_{j}(x,\widehat{y}_{<j},y)\leq\bar{C}$, and $l(y,\widehat{y}^{h}_{j})\leq\bar{l}<\infty$. If $\psi$ is $\psi_{\text{ce}}$ or $\psi_{\text{mae}}$, then with probability $1-\delta$, for a fixed $h$, the excess risk of with respect to ${\cal L}^{\text{OneTime}}$ holds for the empirical risk minimizer $\widehat{r}_{n}$ with respect to ${\cal L}^{\psi}$
@@
where $\tilde{\Gamma}$$\mathbb{R}^{+}\to\mathbb{R}^{+}$, $q(L)=2(|{\cal J}|-1)(\bar{C}+\bar{l})$, $\rho$ is the lipschitz norm of $\psi$, and ${\cal A}_{\psi}\left({\cal F},h\right)={\cal R}^{*}_{\psi}\left({\cal F},h\right)-{\cal R}^{*}_{\psi}(h)$ is the approximation error.
@@
We know that $\psi_{\text{ce}},\psi_{\text{mae}}$ are upper bounded by $u_{\psi_{\text{ce}}}(M),u_{\psi_{\text{mae}}}(M)$ repsectively. We also know that the lipschitz norm for $\psi_{\text{ce}}$ $\rho_{\text{ce}}$ is $\sqrt{2}$ and the lipschitz norm for $\psi_{\text{mae}}$ $\rho_{\text{mae}}$ is 22
@@
${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{ce}}\left({\cal R}_{\text{multi }\psi_{\text{ce}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{ce}}}\right)$ from Theorem 3.1 of Mao et al. [2023] where $\Gamma_{\text{ce}}^{-1}(z)=\frac{1+z}{2}\log(1+z)+\frac{1-z}{2}\log(1-z)$. Similarly, ${\cal R}_{\text{multi }0-1}(f)-{\cal R}^{*}_{\text{multi }0-1}\leq\Gamma_{\text{mae}}\left({\cal R}_{\text{multi }\psi_{\text{mae}}}(f)-{\cal R}^{*}_{\text{multi }\psi_{\text{mae}}}\right)$ where $\Gamma_{\text{mae}}^{-1}(z)=\frac{z}{L+1}$. Both $\Gamma_{\text{ce}}^{-1}$ and $\Gamma_{\text{mae}}^{-1}$ are non-decreasing convex functions making $\Gamma_{\text{ce}}$ and $\Gamma_{\text{mae}}$ non-decreasing concave functions on $\mathbb{R}^{+}$ domain.
@@ Appendix C Baseline Details
For XSUM and TSP, let $p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ be the predictive distribution of the $j^{\text{th}}$ token conditioned on the leftward context $\widehat{y}_{<j}$ and input $x$. The confidence measure would then be $\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ and the predictive entropy would be $H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)=-\sum_{v\in{\cal V}}p_{h}(v\mid\widehat{y}_{<j},x)\log p_{h}(v\mid\widehat{y}_{<j},x)$. For MWP, if we assume that there are dropout layers in the predictor network, we can use the Monte Carlo Variance Gal and Ghahramani [2016] or $\mathrm{Var}_{h}(v\mid y_{<j},x)$ across $T$ inference passes as a confidence measure.
For a given threshold $\tau$, the whole deferral decision rule is determined by evaluating $r(x)>\tau$. The baselines offer various ways of modeling $r$.
  • ChowSum - For TSP and XSUM, $r_{\text{ChowSum}}(x)=-\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ and for MWP, $r_{\text{ChowSum}}(x)=\sum_{j=1}^{L}\mathrm{Var}_{h}(v\mid y_{<j},x)$
  • ChowMean - For TSP and XSUM, $r_{\text{ChowMean}}(x)=-\frac{1}{L}\sum_{j=1}^{L}\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ and for MWP, $r_{\text{ChowSum}}(x)=\frac{1}{L}\sum_{j=1}^{L}\mathrm{Var}_{h}(v\mid y_{<j},x)$
  • ChowQuantile at the level $\alpha$ - For TSP and XSUM, $r_{\text{ChowQuantile}}^{\alpha}(x)=\text{Quantile}_{\alpha}\left(\left\{-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)\right\}_{j=1}^{L}\right)$ and for MWP, $r_{\text{ChowSum}}(x)=\text{Quantile}_{\alpha}\left(\left\{\mathrm{Var}_{h}(v\mid y_{<j},x)\right\}_{j=1}^{L}\right)$
  • WholeModelEmbed - $r(x)$ is trained just like WholeModelScore but the inputs for the rejector are the inputs to the predictor along with the confidence scores.
Similarly, for token-wise deferral, if $r_{j}(x,\widehat{y}_{<j})$ exceeds threshold $\tau$, the expert is called to predict the next token. $r_{j}$ can be modeling in the following ways:
  • TokenwiseScore: For TSP and XSUM, $r_{j}^{\text{score}}(x,\widehat{y}_{<j})=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ and for MSP, $r_{j}^{\text{softmax}}(x,\widehat{y}_{<j})=\mathrm{Var}_{h}(v\mid y_{<j},x)$
  • TokenwiseEntropy: $r_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$. This only applies to XSUM and TSP.
Since $r(x)=\operatorname*{arg\,max}_{j\in[1,L+1]}g_{j}(x)$ for one-time deferral, we consider the following the scoring functions $g_{j}$:
  • OneTimeScore: For TSP and XSUM, $g_{j}^{\text{score}}(x)=-\log p_{h}(\widehat{y}_{j}\mid\widehat{y}_{<j},x)$ and for MSP, $g_{j}^{\text{score}}(x)=\mathrm{Var}_{h}(v\mid y_{<j},x)$
  • OneTimeEntropy: $g_{j}^{\text{entropy}}(x,\widehat{y}_{<j})=H(\widehat{Y}_{j}\mid\widehat{y}_{<j},x)$. This only applies to XSUM and TSP.